# algorithmic

### v3

@@ Appendix A Proof of Theorem 1
and, therefore, $R_{f}(N)\leq\frac{\rho}{2}\left\lVert\nu^{\star}\right\rVert^{2}$ $+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$. Because of Assumption A2, $\frac{1}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}=\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$, and taking into account Assumptions A1.b and A1.c we get $R_{f}(N)\leq\frac{\rho D_{\nu}}{2}+\frac{NG_{f}^{2}}{2\alpha\eta}+\eta(D_{x}^{2}+M_{k_{n}})$. Setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$, we get the sublinear regret bound $R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$. Considering now $R_{c}(N)$, we can rearrange (14) and consider Assumption A1.d, $\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2F}{\rho}+\frac{1}{\rho^{2}}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})+(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})++\frac{\eta}{\rho}\left(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x_{k+1}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}\right).$ Summing from 0 to $N-1$, we get:
$R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2FN}{\rho}+\left\lVert\nu^{\star}\right\rVert^{2}+\frac{\eta}{\rho}\bigg{(}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\bigg{)}.$

### v4

@@ Appendix A Proof of Theorem 1
and, therefore, $R_{f}(N)\leq\frac{\rho}{2}\left\lVert\nu^{\star}\right\rVert^{2}$ $\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$. Because of Assumption A2, $\frac{1}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}=\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$, and taking into account Assumptions A1.b and A1.c we get $R_{f}(N)\leq\frac{\rho D_{\nu}}{2}+\frac{NG_{f}^{2}}{2\alpha\eta}+\eta(D_{x}^{2}+M_{k_{n}})$. Setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$, we get the sublinear regret bound $R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$. Considering now $R_{c}(N)$, we can rearrange (14) and consider Assumption A1.d, $\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2F}{\rho}+\frac{1}{\rho^{2}}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})+(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})++\frac{\eta}{\rho}\left(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x_{k+1}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}\right).$ Summing from 0 to N1N-1, we get:
$R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2FN}{\rho}+\left\lVert\nu^{\star}\right\rVert^{2}$
[…]